OpenStudy (anonymous):

16-2 (8w-7)=3w-18(2-w)

OpenStudy (aroub):

do u know how distribute?

OpenStudy (anonymous):

yes!

OpenStudy (aroub):

so do it and ill check it

OpenStudy (anonymous):

ok

OpenStudy (aroub):

so did u know how to solve it ?

OpenStudy (anonymous):

-2*(8*w-7+16 =3*w-18(2-w)

OpenStudy (anonymous):

taking awhile witing on paper aswell!!

OpenStudy (aroub):

yeah! ofcourse wait give a min

OpenStudy (aroub):

do youhave to find w ?

OpenStudy (anonymous):

yes w=what!

OpenStudy (aroub):

wait please

OpenStudy (aroub):

-37w=-66

OpenStudy (aroub):

is it suppose to be like that?

OpenStudy (anonymous):

w=66/37

OpenStudy (aroub):

yes

OpenStudy (aroub):

did u get it like that?

OpenStudy (anonymous):

yes!

OpenStudy (aroub):

ok :)

OpenStudy (anonymous):

thanks for the help!

OpenStudy (aroub):

anytime :)